OpenStudy (anonymous):

If f(x) = 2x2 + 3x and g(x) = x – 2, what is (f + g)(2)? a. 16 b. 14 c. 12 d. 10 e.8 helpp

OpenStudy (jackellyn):

Since it is asking f+g, you need to add the two expressions together. After simplifying, substitute the x for 2 for your final answer.

OpenStudy (anonymous):

its f(x)=2^2+3x and g(x)=x-2 what is (f+g)(2) so what i got it would equal 22

OpenStudy (jackellyn):

f+g=2x^2+3x+(x-2) Simplify, then substitute the x with 2.

OpenStudy (anonymous):

i got 21

OpenStudy (usukidoll):

the question is asking for (f+g)(2) so let's figure out the f(x) + g(x) portion first. so we have \[f(x)= 2x^2 +3x\] and \[g(x) = x-2 \] so what is f(x)+g(x)

OpenStudy (usukidoll):

then after obtaining f(x)+g(x) we combine like terms finally evaluate f(x)+g(x) when x = 2

OpenStudy (anonymous):

kinda confused still lol sorry

OpenStudy (usukidoll):

let's take it one step at a time by first understanding that (f+g)(2) means that we need to find f+g and then evaluating the result of f+g when x = 2

OpenStudy (anonymous):

okay

OpenStudy (usukidoll):

so we just have to add because f(x) + g(x) means the f function + g function so given \[f(x) = 2x^2+3x \] and \[.g(x) = x-2 \] what is f(x) +g(x) the notation I'm using is the same as f+g

OpenStudy (anonymous):

okay

OpenStudy (usukidoll):

place f(x) here + place g(x) here

OpenStudy (usukidoll):

it should look like \[2x^2+3x+x-2 \] now we combine like terms only for the x portion so what's 3x+x ?

OpenStudy (anonymous):

4x

OpenStudy (usukidoll):

\[2x^2+4x-2\] ok so now we evaluate this function when x = 2. so place x =2 inside of the x of that function ex. \[2(2)^2+4(2)-2\]

OpenStudy (zale101):

(f+g)(2)=f(2)+g(2)

OpenStudy (anonymous):

23?

OpenStudy (usukidoll):

no......

OpenStudy (usukidoll):

what's (2)^2 since we have second power it means to write 2 two times (2)^2 -> (2)(2) what's (2)(2)?

OpenStudy (usukidoll):

(2)(2) notation is the same as 2 x 2 notation

OpenStudy (anonymous):

i mean 14?

OpenStudy (usukidoll):

\[2(2)^2+4(2)-2 \] \[2(4)+4(2)-2 \rightarrow 8+8-2 \rightarrow 16-2=14\]

OpenStudy (usukidoll):

yes it's 14

OpenStudy (anonymous):

thx srry im very slow @learning